c) Assume one of the cluster nodes fails every 100 hours. Other parts of the cluster never fail. Steps 1 through 3 take two hours. On average, the mean time for step 4 is two hours. What is the availability of the cluster? What is the yearly failure cost if each one-hour downtime costs £2000? [10 marks] d) Following Part (b), assume that the cluster now has much increased availability support. Upon a node failure, its workload automatically fails over to other nodes. The failover time is only six minutes. Meanwhile, the cluster has hot swap capability: The faulty node is taken off the cluster, repaired, replugged, and rebooted, and it rejoins the cluster, all without impacting the rest of the cluster. What is the availability of this ideal cluster, and what is the yearly failure cost?
Answer #1

The mean time to failure is = 100 hours given in the question

Mean time to repair is 2+2 = 4 hrs

The availability of this cluster is = 100/(100+4)= 0.9615

This corresponds to a loss of 365*24*(1-0.9615) = 337.26 hrs per year

which means the cost will be 337.26*2000=674520

d)

Availability= MTTF/(MTTF+MTTR)

The cluster's mean time to failure is still 100 hrs but the mean time to repair reduces to 6 mins

So we can see that the availability becomes 100/(100+0.1)=0.999

This corresponds to the loss of 365*24*0.001 = 8.76 hrs per year and 17520 as its cost

Do give a thumbs up and in case there are doubts leave a comment.